OpenStudy (anonymous):

Triple Integration: (Someone to check my approach please) Find the volume of the solid bounded by the cylinder x^2+y^2=4 and the planes y+z=4 and z=0 (i.e. xy plane) My approach is below..

OpenStudy (anonymous):

\[V=\int\limits_{0}^{6}\int\limits_{-2}^{2}\int\limits_{-2}^{2} 1 dxdydz\]

OpenStudy (anonymous):

And I ended up with 93 units cubed

OpenStudy (anonymous):

it's been a bit since i've done these.. but I'm pretty sure you will need to use cylindrical coordinates somewhere there. i.e. you're going to need r and theta

OpenStudy (anonymous):

Yes i also thought so, but it confuses me so much so I tried it in cartesian form...

OpenStudy (anonymous):

Yeah I know what you mean I'm currently fighting with my brain to bring back this stuff aha. well you know x^2+y^2=4 right? that also means r^2=4 or r = 2 so 0<=r<=2 and its a full circle so 0<=theta<=2pi

OpenStudy (anonymous):

yep thats great thanks...except what difference will the other planes make on these limits...thats why i was cheeky and tried to keep it in cartesian....

OpenStudy (anonymous):

then maybe.... z =0 and z=4-y so maybe you can use those as limits?

OpenStudy (anonymous):

and y would be r*sin(theta)

OpenStudy (anonymous):

Oh yes excellent point given that polar would require dr dtheta dz thanks!

OpenStudy (anonymous):

Ill give that a go and compare with my cartesian approach!

OpenStudy (anonymous):

alright good luck!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

hmm do you think i should do it in the order dzdthetadr given that otherwise ill end up with a theta in the final answer?

OpenStudy (anonymous):

theta should be your last bound, z should be your first i think.

OpenStudy (anonymous):

ao dzdrdtheta?

OpenStudy (anonymous):

yeah. and remember when you're doing polar coordinates for that circle its not just dr, you add another r so r*dr dz*r*dr*dtheta

OpenStudy (anonymous):

excellent point thanks!

OpenStudy (anonymous):

okay- I got 200/3 pi hopefully I havent made any mistakes!

OpenStudy (anonymous):

as in 200pi/3

OpenStudy (anonymous):

do you have a way to check your answer?

OpenStudy (anonymous):

um..i wonder if wolfram alpha can do it?

OpenStudy (anonymous):

I wouldnt know what to type in thought...

OpenStudy (anonymous):

though*

OpenStudy (anonymous):

aha oh probably, yeah that's the battle with wolfram.

OpenStudy (anonymous):

I'll have a go...

OpenStudy (anonymous):

was this the integral you drew?

OpenStudy (anonymous):